OpenStudy (meilendurcer):

0.25 L = __mL A. 2.5 B. 25 C. 250 D. 2,500

OpenStudy (anonymous):

C.250

OpenStudy (meilendurcer):

are you sure?

OpenStudy (meilendurcer):

@ktnguyen1 ?

hero (hero):

.25L = xmL Divide both sides by L to get .25 = xm Since milli = 1/1000 re-write as: 25/100 = x/1000 Multiply both sides by 1000: 25*1000/100 = x 25000/100 = x 250 = x

OpenStudy (meilendurcer):

thanks